Whirly Corporation's contribution format income statement for the most recent month is shown below:


TotalPer Unit 
Sales (7,900 units) $ 268,600 $ 34.00
Variable expenses 150,100 19.00
Contribution margin118,500$ 15.00
Fixed expenses54,700
Net operating income$ 63,800

Required:

(Consider each case independently):

1. What would be the revised net operating income per month if the sales volume increases by 70 units?

2. What would be the revised net operating income per month if the sales volume decreases by 70 units?

3. What would be the revised net operating income per month if the sales volume is 6,900 units?



1.Revised net operating income
2.Revised net operating income
3.Revised net operating income
Answer #1

1. Revised net operating income if sales volume increase by 70 units

Sales Volume = 7,900 + 70 = 7,970 units

Particulars Amount
Sales (7,970 units * $34) $270,980
Less: Variable Expense (7,970 units * $19) ($151,430)
Contribution Margin $119,550
Less: Fixed Expense ($54,700)
Net Operating Income $64,850

2. Revised net operating income if sales volume decrease by 70 units

Sales Volume = 7,900 - 70 = 7,830units

Particulars Amount
Sales (7,830 units * $34) $266,220
Less: Variable Expense (7,830 units * $19) ($148,770)
Contribution Margin $117,450
Less: Fixed Expense ($54,700)
Net Operating Income $62,750

3. Revised net operating income if sales volume is 6,900 units

Particulars Amount
Sales (6,900 units * $34) $234,600
Less: Variable Expense (6,900 units * $19) ($131,100)
Contribution Margin $103,500
Less: Fixed Expense ($54,700)
Net Operating Income $48,800
1 Revised Net Operating Income $64,850
2 Revised Net Operating Income $62,750
3 Revised Net Operating Income $48,800

> yo thank you so much for this. i got a similar question like this on my accounting homework and i used this to figure out how to solve the problems

Nana Wordie Fri, Oct 8, 2021 10:34 PM